OpenStudy (anonymous):

Find the tangent of the line to the curve at the given point. cos(x-y)+siny=sqrt(2) (pi/2,pi/4) tell me the steps? please.

OpenStudy (anonymous):

more implicit differentiation and lots of chain rule :D

OpenStudy (anonymous):

\[{\frac {\sin \left( x-y \right) }{\sin \left( x-y \right) +\cos \left( y \right) }} \]

OpenStudy (anonymous):

sin(x-y)/(sin(x-y)+cos(y))

OpenStudy (anonymous):

your m will be 1/2 if you evaluate it at (pi/2,pi/4) then you use y=mx+b plug in y m and x and solve for b